Ok, my algebra teacher gave us this worksheet today and he explained it, and I know what I'm trying to do, but the answer I'm getting dosen't match the equations. We have to write out our own word problem and solve for what we indicated in the problem, like this one, "A theatre sold 200 tickets to a play. The adult tickets cost $3.00 and the student tickets cost $2.50, the number of adult tickets sold was 6 more than 2 times the number of student tickets sold. How many adult tickets were sold?" I f yall can help me I'd really appreciate it. Thank Yall!
Answer by ankor@dixie-net.com(22740)

"A theatre sold 200 tickets to a play. The adult tickets cost $3.00 and the student tickets cost $2.50, the number of adult tickets sold was 6 more than 2 times the number of student tickets sold. How many adult tickets were sold?"
:
This problem seems to be incomplete. You gave the price of the tickets, so it would follow that the problem would include the total $$ received.
:
Using what you have, let a = number of adult tickets and s = number of students
a + s = 200
:
You also say:
"adult tickets sold was 6 more than 2 times the number of student tickets:
a = 2s + 6
:
Substitute (2s + 6) for a in the 1st equation:
(2s + 6) + s = 200
3s = 200 - 6
3s = 194
s = 194/3
s = 64.67 students, fractional students mean there is another flaw in your problem. Perhaps this method will help you